OpenStudy (anonymous):

Solve by the multiplication with add or subtract method 4x - y = -14 3x + 2y = -16

OpenStudy (anonymous):

first multiply the first equation by 2 [4x-y=-14]*2 8x-2y=-28 add the product to the second equation 3x+2y=-16 + 8x-2y=-28 --------------------- 11x=-44 x=-44/11 x=-4 plug in the value x in one of the equations 4(-4)-y=-14 -16-y=-14 y=-16+14 y=-2

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Anytime :)